1) I) The manager of a furniture manufacturing plant hopes to achieve a better allocation of...

1)

I)

The manager of a furniture manufacturing plant hopes to achieve a better allocation of inventory control efforts by adopting an ABC approach to inventory control. Given the monthly usages listed below, classify the items in A, B, and C categories according to dollar usage.

Annual Usage

Unit Cost

Item Type

in thousands

in Dollars

Item 1

10

$2.50

Item 2

30

$1.25

Item 3

20

$16.00

Item 4

10

$0.75

Item 5

20

$3.00

Item 6

33

$2.50

Item 7

60

$14.25

Item 8

22

$10.00

Item 9

100

$2.75

Item 10

500

$4.25


Using the above data, Item 6 will be classified as:

  • A) A item
  • B) B item
  • C) C item
  • D)  None of the above

ii) Can you show the formulas and steps for this question
:

A project network is given by the following table. The minimum time required to complete the project is:

Activity

Immediate Predecessors

Activity Time

A

-

   5 days

B

A

   9 days

C

A

   8 days

D

B, C

   7 days

E

C

   4 days

F

E

   3 days

G

D, F

   3 days

  • A.) 23 days
  • B.) 24 days
  • C) 25 days
  • D.)39 days
  • E.) none of the above
Answer #1

i) C) C item

Perform the following operations

Annual Usage in Dollars = Annual Usage in thousands * Unit cost in dollars

Sort the items in descending order of Annual Usage in Dollars

Annual Dollar Usage % of Total = Annual Usage in Dollars / Total

Cumulative % = Running SUM of Annual Dollar Usage % of Total

Item Type Annual Usage in thousands Unit Cost in Dollars Annual Usage in Dollars (1000) Annual Dollar Usage % of Total Cumulative %
Item 10 500 4.25 2125 53.0% 53.0%
Item 7 60 14.25 855 21.3% 74.4%
Item 3 20 16 320 8.0% 82.3%
Item 9 100 2.75 275 6.9% 89.2%
Item 8 22 10 220 5.5% 94.7%
Item 6 33 2.5 82.5 2.1% 96.8%
Item 5 20 3 60 1.5% 98.3%
Item 2 30 1.25 37.5 0.9% 99.2%
Item 1 10 2.5 25 0.6% 99.8%
Item 4 10 0.75 7.5 0.2% 100.0%
Total 4007.5

Items comprising the last 5% of Total Annual Dollar Usage are classified as C. Item 6 is in the last 5%, hence classified as C item

----------------------------------------------------------------------------------------------------

ii) B) 24 days

List alternative paths in the network and their duration. Duration of each path is equal to the sum of activity times on that path.

Path Duration
A-B-D-G 5+9+7+3 = 24
A-C-D-G 5+8+7+3 = 23
A-C-E-F-G 5+8+4+3+3 = 23

Path of longest duration is the critical path.

Critical path is A-B-D-G

Minimum time required to complete the project is equal to the time of the critical path, which is 24 days